OpenStudy (anonymous):

Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden?

OpenStudy (anonymous):

\[208=(10+x)*(13+x)\]\[208=x^{2}+23x+130\]\[x^{2}+23x-78=0\]\[(x-3)(x+26)=0\]\[X=3; x=-26\] The longer side will be: \[13+3=16m\] CHeck \[(13+3)*(10+3)=208\]\[208=208\] CHECK

OpenStudy (anonymous):

16? Thanks I got it right :)